Sol Summary – February 6th

Sol: Sol #2
Summary Title: Training and raining – the rain, the rain, (and we have only just begun our work).
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Activity Summary: Today we ate breakfast together and were ready for our training EVA’S well ahead of time. We were all there to help each other with the necessary preparation tasks of the EVA – to dress efficiently and comfortably in the space suits, to be completely at ease with the communication equipment as well as to be able to drive the electric rovers to our destination and move about in the field whilst wearing the space suits. We split into two groups and two separate EVAS to Marble Ritual. Both excursions were very successful and well time-managed despite a costume malfunction (mine) and drizzling rain which has now become intense. The rain is cramping our EVA ambitions. After the second training EVA, today we met at four to discuss and share our research goals and to determine the science directions for the next few days. We discussed everyone’s research and talked about what each of us needs next, and what we can do to assist. We talked about the future, about Ladakh and plans to build an analogue research station, an observatory and a green hab there. We talked about our initial impressions of MDRS. Our microbiologist Daniel Loy and geologist Bharti Sharma contacted Science Director and MSA president geologist Dr. Jon Clarke for more advice about where to go for gypsum sites within the range of the rovers. My feeling is that overall the crew is working well together. Our health is good. We are fit and sleeping and eating well. Fingers crossed that the rain will cease and we can go out in the field.
Look Ahead Plan: An EVA tomorrow for four crew to do slope analysis and find gypsum samples. A day at home in the hab for the rest of the crew with work here to do.
Anomalies in work: Nothing to report.
Weather: Cool start to the morning with 30 degrees F at seven am and humidity at 81%. High cloud cover all day and a slight wind with the temperature rising to 49F. Rain began at about 2.30 pm with drizzle then pouring, temperatures dropping quickly and high winds expected.
Crew Physical Status: Nominal
EVA: Training EVA #1 today to Marble Ritual at 10.30am with Mehanz, Rajvi, Clare and Aditya. Training EVA # 2 to Marble Ritual at 1.45 with Bharti, Daniel and me, Annalea.
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, Astronomy Report
EVA Request(s): EVA with Bharti, Daniel, Aditya and Rajvi. Leaving at 2pm to Cowboy Corner in search of gypsum samples as well as slope analysis.
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site